40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6520     Entry time: Wed Apr 11 16:33:16 2012
Author: Jenne 
Type: Bureaucracy 
Category: General 
Subject: 40m Meeting Action Items 

Action Items from Last Week:

(non-existant)

-----------------------------------------------------------------------

Action Items this Week and LEAD PERSON:

Assemble and ship 4 TTs from LHO - SURESH

Prepare electronics for TTs (coil drivers) - JAMIE

In-air TT testing to confirm we can control / move TTs before we vent (starting in ~2 weeks) - SURESH

Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH

OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN

Static-only OAF noise budget (Adaptive noise budget as next step) - DEN

Black glass: big baffle pieces to clean&bake, get small pieces from Bob, put into baskets, make new basket for 1" pieces, get to clean&bake - KOJI

IPPOS beam measurement - SURESH with JENNE

AS beam measurement (if beam is bright enough) - SURESH and JENNE

Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE

Summary of IFO questions, measurements to take, and game plan - JENNE

Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH

ELOG V3.1.3-